# Hierarchical Correlation Clustering and Tree Preserving Embedding

### v2

@@ 4.2 Embedding the HCC-based distances
    We center $\mathbf{X}$ by
    $\mathbf{B}\leftarrow-\frac{1}{2}(\mathbf{I}_{n}-\frac{1}{n}\mathbf{e}_{n}\mathbf{e}_{n}^{\texttt{T}})\mathbf{X}(\mathbf{I}_{n}-\frac{1}{n}\mathbf{e}_{n}\mathbf{e}_{n}^{\texttt{T}}),$ (25)
    where $\mathbf{e}_{n}$ is an $n-$dimensional constant vector of $1$’s and $\mathbf{I}_{n}$ is an identity matrix of size $n\times n$. With this centering, the sum of both the rows and columns in matrix $\mathbf{B}$ becomes zero.
@@
    $\mathbf{B}=\mathbf{Y}\mathbf{Z}\mathbf{Y}^{\texttt{T}},$ (26)
@@
    $\mathbf{A}=\mathbf{Y}_{l}(\mathbf{Z}_{l})^{1/2},$ (27)
In the embedded space, the new dimensions are arranged based on their corresponding eigenvalues. One can opt to select only the most significant ones, rather than utilizing all of them. Therefore, computing such an embedding also offers the benefit of feature selection. We also note that many clustering methods can be written in matrix factorization form via for example spectral $K$-means [25]. This induces an embedding and hence a set of relevant features. However, for general positive/negative similarity matrices no exact embedding is feasible due to violating positive semidefiniteness. The method that we described here provides a solution to this challenge, i.e., enables us to extract features when the base pairwise similarities are positive and negative.
5 Correlation Clustering and Minimax Dissimilarities
Finally, we study minimax dissimilarities for correlation clustering, a graph-based method that corresponds to constructing a hierarchical clustering. Given graph $\mathcal{G}(\mathbf{O},\mathbf{D})$, the minimax (MM) dissimilarity between $i$ and $j$ is defined as
$\mathbf{D}_{i,j}^{MM}$ $=$ $\min_{p\in\mathcal{P}_{ij}(\mathcal{G})}\max_{1\leq l\leq|p|-1}\mathbf{D}_{p(l)p(l+1)},$ (28)
Minimax dissimilarities enable a clustering algorithm to capture the underlying patterns and manifolds in an unsupervised and nonparametric way via extracting the transitive relations [42, 14]. For example, if object $i$ is similar to object $j$, $j$ is similar to $k$, and $k$ is similar to $l$, then the minimax dissimilarity between $i$ and $l$ will be small, even though their direct dissimilarity might be large. The reason is that minimax dissimilarity finds the connectivity path $i\to j\to\cdots\to k\to l$ and connects $i$ and $l$ via this path. This property is helpful in finding elongated clusters and manifolds of arbitrary shapes in an unsupervised way.
Lemma 1
All the proofs are in Appendix A. Hence, given a dissimilarity matrix $\mathbf{D}$, one can subtract $\alpha:=min(\mathbf{D})$ from all the elements to obtain $\mathbf{D}^{\alpha}$. Then, the minimax dissimilarities can be computed from $\mathcal{G}^{\alpha}(\mathbf{O},\mathbf{D}^{\alpha})$. After computing the minimax dissimilarities from $\mathcal{G}^{\alpha}$, we may add $\alpha$ to all the pairwise minimax dissimilarities. We can obtain the minimax similarities $\mathbf{S}^{MM}_{ij}$ via $\mathbf{S}^{MM}_{ij}=-\mathbf{D}^{MM}_{ij}$, if needed.
Theorem 1
As mentioned, correlation clustering on arbitrary similarity matrix $\mathbf{S}$ is NP-hard [9, 23] and even is APX-hard [23]. Therefore, using minimax (dis)similarities with correlation clustering not only helps for extracting elongated complex patterns, but also yields a significant reduction in the computational complexity, i.e., from NP-hardness to a polynomial runtime.
Among the approximate algorithms proposed for correlation clustering on complete graphs with discrete weights, the method in [6] provides a 3-factor approximation. with a randomly selected unclustered object in the graph, this method greedily finds the object’s positive neighbors (those with similarity $+1$) to form a new cluster. Then, it repeats this procedure for the remaining objects. One can conclude that in the optimal solution of correlation clustering on graph $\mathcal{G}(\mathbf{O},\mathbf{S}^{MM})$, exactly only the positive neighbors of an object will be in the same cluster as the object is, i.e., interestingly the 3-factor approximation algorithm in [6] becomes exact when applied to $\mathcal{G}(\mathbf{O},\mathbf{S}^{MM})$ (Theorem 2).
Theorem 2
6 Experiments
In this section, we describe our experimental results on various datasets. We compare our methods with single linkage (SL), complete linkage (CL) and average linkage (AL).444 Some criteria, e.g. centroid, median and Ward compute a representative for each cluster and then compute the inter-cluster dissimilarities by the distances between the representatives. Computing such representatives might not be feasible for possibly negative pairwise dissimilarities. Thus, we do not consider them. As mentioned, there are several improvements over these basic methods. However, such contributions are orthogonal to our contribution. Moreover, it is unclear how such improvements can be extended to the dissimilarities that can be both positive and negative. Thus, we limit our baselines to these methods which as mentioned in Proposition 1, are applicable to such data.
(a) [Breast Tissue]
(b) Cardio
(c) Image Segmentation
(d) ISOLET
(e) Leaf
(f) One-Hundred Plant
Figure 2: MI score of different hierarchical clustering methods applied to UCI datasets, where x-axis shows the parameter $\eta$. HCC significantly yields the best results.
@@ 6 Experiments
Table 1: Performance of different tree preserving embedding methods on UCI datasets applied with GMM. The embeddings often improve the results, in particular HCC+GMM when the task is difficult.
Breast Tissue Cardiotocography Image Segm. ISOLET Leaf One-Hun. Plant
method MI Rand MI Rand MI Rand MI Rand MI Rand MI Rand
SL 0.006 0.008 0.007 0.007 0.008 0.001 0.009 0.016 0.008 0.003 0.015 0.020
SL+GMM 0.093 0.077 0.120 0.135 0.239 0.250 0.192 0.174 0.155 0.161 0.083 0.077
CL 0.227 0.166 0.077 0.056 0.187 0.125 0.057 0.017 0.081 0.038 0.029 0.008
CL+GMM 0.251 0.171 0.081 0.060 0.201 0.154 0.061 0.043 0.140 0.129 0.054 0.049
AL 0.542 0.519 0.391 0.479 0.518 0.495 0.257 0.165 0.181 0.106 0.066 0.023
AL+GMM 0.550 0.513 0.422 0. 463 0.522 0.501 0.240 0.179 0.152 0.143 0.081 0.065
HCC 0.903 0.900 0.987 0.994 0.945 0.943 0.938 0.918 0.429 0.373 0.159 0.104
HCC+GMM 0.914 0.911 0.979 0.974 0.960 0.966 0.941 0.917 0.462 0.401 0.183 0.217
6.1 HCC on UCI data
We first investigate the hierarchical correlation clustering on the following six UCI datasets chosen mainly from the image/shape domain [41].
@@ 6.1 HCC on UCI data
    Breast Tissue: includes electrical impedance measurements of freshly excised $106$ tissue samples from the breast. The number of clusters is $6$.
@@
    Cardiotocography: contains $2126$ measurements of fetal heart rate and uterine contraction features on cardiotocograms in $10$ clusters.
  3. 3.
    Image Segmentation: contains $2310$ samples from images of $7$ outdoor clusters.
  4. 4.
    ISOLET: $7797$ samples consisting of spoken attributes of different letters ($26$ clusters).
  5. 5.
    Leaf: $340$ images of leaf specimens originating from $40$ different plant species (clusters) each described by $16$ attributes.
  6. 6.
    One-Hundred Plant: $1600$ samples of leafs each described by $64$ features, from in total $100$ types (clusters).
The ground-truth labels are shown by $\mathbf{c}^{*}_{i}$. We assume an oracle reveals the pairwise similarities $\mathbf{S}$ according to the (flip) noise parameter $\eta$. If $\mathbf{c}^{*}_{i}=\mathbf{c}^{*}_{j}$ then $\mathbf{S}_{i,j}=\mathcal{U}(0,1)$ with probability $1-\eta$ and $\mathbf{S}_{i,j}=\mathcal{U}(-1,0)$ with probability $\eta$. If $\mathbf{c}^{*}_{i}\neq\mathbf{c}^{*}_{j}$ then $\mathbf{S}_{i,j}=\mathcal{U}(-1,0)$ with probability $1-\eta$ and $\mathbf{S}_{i,j}=\mathcal{U}(0,1)$ with probability $\eta$. The function $\mathcal{U}(.,.)$ returns a uniform random number within the specified range. For each $\eta$ we repeat the experiments $20$ times and report the average results. This setup provides a systematic approach to study the robustness of various methods to noise.
Figure 2 shows the results for different datasets as a function of the noise level $\eta$ w.r.t. MI. Rand scores shown in Figure 3 exhibits a consistent behavior. We observe that among different methods, HCC performs significantly better and produces more robust clustering w.r.t. the noise parameter $\eta$. The results on Leaf and One-Hundred Plant are worse with all the methods. The reason is that these datasets are complex, having many clusters (respectively, $40$ and $100$ clusters) and fairly a small number of objects per cluster.
6.2 Tree preserving embedding of UCI data
In the following, we investigate tree preserving embedding and feature extraction. After computing the embeddings from different hierarchical clustering methods, we apply Gaussian Mixture Model (GMM) to the extracted features and evaluate the final clustering solution. This kind of embedding enables us to apply methods such as GMM to positive and negative similarities, a task that was not possible before. Since the extracted features appear in the form of vectors, thus, the final clustering method is not limited to GMM, and in principle, any numerical clustering is applicable.
Table 1 shows the results for different UCI datasets. We observe that the embeddings obtained by HCC (i.e., ‘HCC+GMM’) yield significantly better results compared to the embeddings from other methods (i.e., ‘SL+GMM’, ‘CL+GMM’ and ‘AL+GMM’). It’s worth noting that the results of embeddings (e.g., ‘HCC+GMM’) typically surpass the results obtained from the hierarchical clustering alone (e.g., ‘HCC’). This observation supports the idea that employing HCC to compute an embedding (for extracting new features) for a clustering method such as GMM may be advantageous, yielding superior results compared to using HCC exclusively for clustering purposes. This verifies why tree preserving embedding can be effective in general.
6.3 Experiments on Fashion-MNIST
Fashion MNIST 1 Fashion MNIST 2
method MI Rand MI Rand
SL 0.322 0.206 0.241 0.196
SL+GMM 0.411 0.335 0.384 0.340
CL 0.403 0.293 0.546 0.379
CL+GMM 0.478 0.426 0.574 0.362
AL 0.464 0.468 0.602 0.534
AL+GMM 0.608 0.551 0.647 0.553
HCC 0.499 0.475 0.666 0.557
HCC+GMM 0.581 0.586 0.693 0.548
@@ 6.4 HCC on 20 newsgroup data
    news1: all the $3901$ documents of the categories ’misc.forsale’, ’rec.motorcycles’, ’talk.politics.mideast’, ’sci.med’ ($48596$ dimensions).
@@
    news2: all the $3743$ documents of the categories ’alt.atheism’, ’comp.sys.mac.hardware’, ’sci.electronics’, ’soc.religion.christian’ ($40735$ dimensions).
@@
    news3: all the $1984$ documents of the categories ’sci.space’, ’soc.religion.christian’ ($30749$ dimensions).
@@
    news4: all the $2877$ documents of the categories ’comp.graphics’, ’rec.sport.baseball’, ’talk.politics.guns’ ($38177$ dimensions).
@@
For each dataset, we compute the TF-IDF vectors of the documents and apply PCA with $50$ principal components. [^5] The results are consistent with different number of components. We obtain the similarity between every two documents via the cosine similarity between their respective PCA vectors, which is a number in $[-1,+1]$. We then investigate the different hierarchical clustering methods. Table 3 shows the results w.r.t. different evaluation criteria. Among the different methods, HCC usually yields the best results, and AL is the second best choice.
@@
In the following, we investigate tree-preserving embedding on these datasets. Table 4 presents the results of tree-preserving embedding compared to hierarchical clustering. Specifically, we investigate the benefits of using hierarchical clustering for feature extraction. We observe, i) employing hierarchical clustering to extract features for a method such as GMM usually yields improving results, and ii) HCC, whether used directly for clustering or for feature extraction, often yields superior results compared to alternatives.
6.5 Experiments with web data
@@
Finally, we investigate HCC for both hierarchical clustering and feature extraction on a web dataset. The dataset consists of $15,000$ webpages collected about topics such as politics, finance, sport, art, entertainment, health, technology, environment, cars and films. Similar to 20 newsgroup datasets, we compute the TF-IDF vectors, apply PCA and then obtain the pairwise cosine similarities between the web pages. Table 5 demonstrates the performance of different methods on this dataset. We observe that, consistent with the previous experiments, both HCC and HCC+GMM yield improving the results compared to the baselines. In addition, using HCC to compute intermediate features for GMM (i.e., HCC+GMM) provides slightly better scores than using HCC to produce the final clusters.
(a) DS2 [40, 59]
(b) DS4 [19]
(c) DS5 [1]
(d) DS6 [1]
6.6 Correlation clustering with minimax dissimilarities
Finally, we investigate the use of minimax dissimilarities with correlation clustering. As mentioned, minimax dissimilarities are in particular useful for extracting elongated and complex patterns in an unsupervised way. Thereby, we apply ‘minimax + correlation clustering’ to a number of datasets with visually elongated and arbitrarily shaped clusters, and compare the results with ‘correlation clustering’ alone. The datasets are shown in Figure 4. For each dataset, we simply construct the K-nearest neighbor graph (using the ordinary Euclidean distance with a typical K such as 3). The edges in the nearest neighborhood receive a positive weight (e.g., +1) and a negative weight (e.g., -1) otherwise. We then apply either ‘minimax + correlation clustering’ or ‘correlation clustering’ to the resultant graph. We observe that for all the datasets, ‘minimax + correlation clustering’ yields the perfect clustering, i.e., MI and Rand are equal to 1. Whereas with ‘correlation clustering’ alone these scores are low, less than 0.2. We acknowledge the existence of other methods, such as DBSCAN [27], capable of achieving perfect clustering on these datasets. However, these methods often involve crucial hyperparameters, the tuning of which can be challenging in unsupervised learning. To our knowledge, in addition to computational and theoretical benefits, ‘minimax + correlation clustering’ stands out as the sole method capable of achieving perfect results on datasets with elongated clusters of arbitrary shapes, and fulfills the following two promises: i) it eliminates the need to fix hyperparameters, a task often intricate in unsupervised learning, and ii) it automatically determines the correct number of clusters without requiring prior fixing.
7 Conclusion
We proposed a new hierarchical clustering method, called HCC, suitable when the (dis)similarities can be positive or negative. This method is consistent with the general algorithmic procedure for agglomerative clustering and only differs in the way the inter-cluster dissimilarity function is defined. We then developed a framework for embedding the HCC dendrograms. Such an embedding provides extracting useful features to apply for example a GMM for clustering positive and negative similarities. In the following, we studied the use of minimax dissimilarities with correlation clustering and showed that it yields reduction in the computational complexity, in addition to a possibility for extracting elongated manifolds. Finally, we demonstrated the effectiveness of our methods on several datasets in different settings.
Acknowledgement
The work of Morteza Haghir Chehreghani was partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation.
Appendix A Proofs
A.1 Proof of Lemma 1
One can show that the pairwise minimax dissimilarities on any arbitrary graph are equivalent to the pairwise minimax dissimilarities on any minimum spanning tree derived from the same graph. The proof is similar to the maximum capacity problem [36]. Thereby, the minimax dissimilarities are obtained by
$=$ $\max_{1\leq l\leq|p_{ij}|-1}\mathbf{D}_{p(l)p(l+1)},$
where $p_{ij}$ indicates the (only) path between $i$ and $j$ on the minimum spanning tree. To obtain the minimax dissimilarities $\mathbf{D}^{MM}_{ij}$, we select the maximal edge weight on the only path between $i$ and $j$ on the minimum spanning tree.
On the other hand, single linkage method and the Kruskal’s minimum spanning tree algorithm are equivalent [31]. Thus, the dendrogram $T$ obtained via single linkage sufficiently contains the pairwise minimax dissimilarities. Now, we need to show that the minimax dissimilarities in Eq. 29 equal the dissimilarities defined in Eq. 20, i.e., $\mathbf{X}_{ij}$ is the largest edge weight on the path between $i$ and $j$ in the hierarchy.
Given $i,j$, let
A.2 Proof of Theorem 2
  1. 1.
  2. 2.
    Whenever there is a positive path between $i$ and $j$, then their minimax similarity $\mathbf{S}^{MM}_{i,j}$ will be necessarily positive. Therefore, when we apply correlation clustering to graph $\mathcal{G}(\mathbf{O},\mathbf{S}^{MM})$, all the intra-cluster similarities of the optimal clusters will be positive. This corresponds to having a positive path between every two objects that are in the same optimal cluster, i.e., they are in the same connected component of $\mathcal{G}(\mathbf{O},\mathbf{S}^{\prime})$ where $S^{\prime}$ is defined as
  3. 3.
  • If two nodes $i$ and $j$ are at two different connected components, then there is no positive path between them either on $\mathcal{G}(\mathbf{O},\mathbf{S}^{\prime})$ or on $\mathcal{G}(\mathbf{O},\mathbf{S}^{MM})$. Thus, they cannot be at the same cluster if we apply correlation clustering on $\mathcal{G}(\mathbf{O},\mathbf{S}^{MM})$.
A.3 Proof of Theorem 3
The approximate algorithm in [6] iteratively picks an unclustered object and its positive neighbors as a new cluster. According to Theorem 1, the optimal solution of correlation clustering applied to $\mathcal{G}(\mathbf{O},\mathbf{S}^{MM})$ corresponds to extracting the connected components of graph $\mathcal{G}(\mathbf{O},\mathbf{S}^{\prime})$, where $S^{\prime}$ is defined in Eq. 31.
Thus it is sufficient to show the two followings.
  1. 1.
  2. 2.

### v4

@@ B.2 Proof of Theorem 2
    Whenever there is a positive path between $i$ and $j$, then their minimax similarity $\mathbf{S}^{MM}_{i,j}$ will be necessarily positive. Therefore, when we apply correlation clustering to graph $\mathcal{G}(\mathbf{O},\mathbf{S}^{MM})$, all the intra-cluster similarities of the optimal clusters will be positive. This corresponds to having a positive path between every two objects that are in the same optimal cluster, i.e., they are in the same connected component of $\mathcal{G}(\mathbf{O},\mathbf{S}^{\prime})$ where $\mathbf{S}^{\prime}$ is defined as
@@
  • If two nodes $i$ and $j$ are at two different connected components, then there is no positive path between them either on $\mathcal{G}(\mathbf{O},\mathbf{S}^{\prime})$ or on $\mathcal{G}(\mathbf{O},\mathbf{S}^{MM})$. Thus, they cannot be in the same cluster if we apply correlation clustering on $\mathcal{G}(\mathbf{O},\mathbf{S}^{MM})$.
@@
B.3 Proof of Theorem 3
The approximate algorithm in [5] iteratively picks an unclustered object and its positive neighbors as a new cluster. According to Theorem 1, the optimal solution of correlation clustering applied to $\mathcal{G}(\mathbf{O},\mathbf{S}^{MM})$ corresponds to extracting the connected components of graph $\mathcal{G}(\mathbf{O},\mathbf{S}^{\prime})$, where $\mathbf{S}^{\prime}$ is defined in Eq. 17.
Thus it is sufficient to show the two followings.
@@ B.3 Proof of Theorem 3
Appendix C Additional Experimental Results
C.1 More results with UCI datasets
In Figure 5, we demonstrate the performance of HCC on different UCI datasets w.r.t. Rand score, and compare the results with other agglomerative methods (the results w.r.t. MI are shown in the main text in Figure 2). We observe that consistent with the MI measure, HCC yields the best scores, even at high noise levels and when the datasets are difficult to cluster.
@@ C.2 HCC on 20 newsgroup data
    news1: the $3901$ documents of the categories ’misc.forsale’, ’rec.motorcycles’, ’talk.politics.mideast’, ’sci.med’ ($48596$ dimensions).
@@
    news2: the $3743$ documents of the categories ’alt.atheism’, ’comp.sys.mac.hardware’, ’sci.electronics’, ’soc.religion.christian’ ($40735$ dimensions).
@@
    news3: the $1984$ documents of the categories ’sci.space’, ’soc.religion.christian’ ($30749$ dimensions).
@@
    news4: the $2877$ documents of the categories ’comp.graphics’, ’rec.sport.baseball’, ’talk.politics.guns’ ($38177$ dimensions).
@@
For each dataset, we compute the TF-IDF vectors of the documents and apply PCA with $50$ principal components. We obtain the similarity between every two documents via the cosine similarity between their respective PCA vectors. As has been discussed in detail in [37], adding a fixed number to all the pairwise similarities can possibly improve the clustering results. [^5] It is suggested in [18] to adaptively shift the pairwise similarities so that the sum of pairwise similarities equals zero for each object in the dataset. Table 3 shows the results for various hierarchical clustering methods w.r.t. different evaluation criteria. Among the different methods, HCC usually yields the best results, and AL is the second best choice.
@@
In the following, we investigate tree-preserving embedding on these datasets. Table 4 presents the results of tree-preserving embedding compared to hierarchical clustering. Specifically, we investigate the benefits of using hierarchical clustering for feature extraction. We observe, i) employing hierarchical clustering to extract features for a method such as GMM usually yields improving the results, and ii) HCC, whether used directly for clustering or for feature extraction, often gives superior results compared to the alternatives.
C.3 Experiments with web data
@@
Finally, we investigate HCC for both hierarchical clustering and feature extraction on a web dataset. The dataset consists of $15,000$ webpages collected about topics such as politics, finance, sport, art, entertainment, health, technology, environment, cars and films. Similar to 20 newsgroup datasets, we compute the TF-IDF vectors, apply PCA and then obtain the pairwise cosine similarities between the web pages. Table 5 demonstrates the performance of different methods on this dataset. We observe that, consistent with the previous experiments, both HCC and HCC+GMM yield improving the results compared to the baselines. In addition, using HCC to compute intermediate features for GMM (i.e., HCC+GMM) results in better scores than using HCC to produce the final clusters.